# Fixed-Time Input-to-State Stability for Singularly Perturbed Systems via Composite Lyapunov Functions

### v1

@@ 4 Fixed-Time Feedback Optimization with Time-varying Costs
$\left\lvert\begin{bmatrix}\tilde{x}(t)\\ \tilde{z}(t)\end{bmatrix}\right\rvert\leq\beta_{\varepsilon}\left(\left\lvert\begin{bmatrix}\tilde{x}(0)\\ \tilde{z}(0)\end{bmatrix}\right\rvert,t\right)+\vartheta_{\varepsilon}(|\varepsilon_{0}|),$
for all $t\geq 0$.
Proof 4.2.
To show FxT ISS, we will first consider a version of (28) in which (28b) is “unperturbed”.
$\dot{{\tilde{x}}}$ $=-\dfrac{\nabla\phi_{\theta}(\tilde{z}+\varphi(\theta))}{|\nabla\phi_{\theta}(\tilde{z}+\varphi(\theta))|^{\xi_{1}}}-\dfrac{\nabla\phi_{\theta}(\tilde{z}+\varphi(\theta))}{|\nabla\phi_{\theta}(\tilde{z}+\varphi(\theta))|^{\xi_{2}}}-\varepsilon_{0}\dfrac{d\varphi(\theta)}{d\theta}\Pi(\theta)$ (31a)
$\varepsilon\dot{{\tilde{z}}}$ $=g(\tilde{x}+\varphi(\theta),\tilde{z}+\varphi(\theta)),$ (31b)
which we will show is FxT ISS with input $\varepsilon_{0}$. The system has quasi steady state $h(\tilde{x})=\tilde{x}$ with the following reduced system:
$\dot{\tilde{x}}$ $=-\dfrac{\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))}{|\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))|^{\xi_{1}}}-\dfrac{\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))}{|\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))|^{\xi_{2}}}-\varepsilon_{0}\dfrac{d\varphi(\theta)}{d\theta}\Pi(\theta).$ (32)
We can let $M=\max_{\theta\in\Theta}|\frac{d\varphi(\theta)}{d\theta}\Pi(\theta)|$. Then, with the FxT ISS Lyapunov function $V(x)=|\tilde{x}|^{2}$ we can compute its Lie derivative along the trajectories of (32) to obtain:
$\dot{V}$ $=-2\dfrac{\tilde{x}^{\top}\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))}{|\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))|^{\xi_{1}}}-2\dfrac{\tilde{x}^{\top}\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))}{|\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))|^{\xi_{2}}}$
$~{}~{}~{}-2\varepsilon_{0}\tilde{x}^{\top}\dfrac{d\varphi(\theta)}{d\theta}\Pi(\theta)$
$\leq-\frac{2}{L}\left(|\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))|^{2-\xi_{1}}+|\nabla\phi_{\theta}({\tilde{x}}+\varphi(\theta))|^{2-\xi_{2}}\right)$
$~{}~{}~{}+2M|\tilde{x}||\varepsilon_{0}|$
$\leq-\frac{2}{L}\left(\kappa^{2-\xi_{1}}|\tilde{x}|^{2-\xi_{1}}+\kappa^{2-\xi_{2}}|\tilde{x}|^{2-\xi_{2}}\right)+\frac{M}{\underline{\alpha}_{R}(c)}|\tilde{x}|^{2}$
$~{}~{}~{}+M\overline{\alpha}_{R}(c)|\varepsilon_{0}|^{2}$
$\leq-\frac{\kappa^{2-\xi_{1}}}{L}V^{1-\frac{1}{2}\xi_{1}}-\frac{\kappa^{2-\xi_{2}}}{L}V^{1-\frac{1}{2}\xi_{2}}+M\overline{\alpha}_{R}(c)|\varepsilon_{0}|^{2},$
where $\underline{\alpha}_{R},\overline{\alpha}_{R}\in\mathcal{K}_{\infty}$ are from Lemma 1 and $c$ satisfies
$\frac{M}{\underline{\alpha}(c)}\leq\min\left(\frac{\kappa^{2-\xi_{1}}}{L},\frac{\kappa^{2-\xi_{2}}}{L}\right).$
From this, we conclude (32) satisfies Assumption 2. Let $W$ be obtained from Assumption 4, where we can assume without loss of generality that $\beta_{1}=1$. Then, we can use (10) to obtain
$\tilde{V}$ $=V^{\frac{1}{2}-\frac{1}{4}\xi_{1}}+V^{\frac{1}{2}-\frac{1}{4}\xi_{2}}$ (33a)
$\tilde{W}$ $=W^{\frac{1}{2}b_{1}}+W^{\frac{1}{2}b_{2}}.$ (33b)
To obtain FxT ISS, it only remains to check the interconnection conditions. We can define $\Phi_{\theta}(x):=\nabla\phi_{\theta}(x+\varphi(\theta))$ and compute $I_{1}$:
$I_{1}$ $=2\tilde{x}^{\top}\biggl{(}\dfrac{\Phi_{\theta}(\tilde{x})}{|\Phi_{\theta}(\tilde{x})|^{\xi_{1}}}+\dfrac{\Phi_{\theta}(\tilde{x})}{|\Phi_{\theta}(\tilde{x})|^{\xi_{2}}}-\dfrac{\Phi_{\theta}(y+\tilde{x})}{|\Phi_{\theta}(y+\tilde{x})|^{\xi_{1}}}$
$~{}~{}~{}-\dfrac{\Phi_{\theta}(y+\tilde{x})}{|\Phi_{\theta}(y+\tilde{x})|^{\xi_{2}}}\biggl{)}.$
Figure 2: The trajectories of the FxT time-varying feedback optimization interconnection with $\varepsilon_{0}=0.7,0.5,0.2$ and $\varepsilon=0.001$. The system is initialized at $x_{0}=[500;333.3]$ and $z_{0}=[690;210]$
where $y=\tilde{z}-\tilde{x}$. We can use (29) and Lemma 3 to obtain the following estimate:
$I_{1}$ $\leq 2^{\xi_{1}+1}L^{1-\xi_{1}}|\tilde{x}||y|^{1-\xi_{1}}$
$~{}~{}~{}+2L^{1-\xi_{2}}K_{\xi_{2}}\left(|\tilde{x}|^{1-\xi_{2}}|y|+|\tilde{x}||y|^{1-\xi_{2}}\right)$
$\leq\dfrac{1}{\underline{\alpha}_{1}(c)}\left(|\tilde{x}|^{2-\xi_{1}}+|\tilde{x}|^{2-\xi_{2}}\right)+\overline{\alpha}_{1}(c)\left(|y|^{2-\xi_{1}}+|y|^{2-\xi_{2}}\right)$
$\leq\dfrac{1}{\underline{\alpha}_{1}(c)}\tilde{V}^{2}+2\overline{\alpha}_{1}(c)\tilde{W}^{2},$
for all $c>0$, where $\underline{\alpha}_{1},\overline{\alpha}_{1}\in\mathcal{K}_{\infty}$. So (11a) holds, and we can then pick $c$ sufficiently large such that (11c) is satisfied. Now we only need to check (11b), which can be verified by using (26) and (29). Due to space limitations, this computation is omitted.
Thus, by Theorems 1 and 2 we know there exists $\varepsilon^{*}$ such that (31) is FxT ISS with input $\varepsilon_{0}$ for $\varepsilon\in(0,\varepsilon^{*})$. Namely, there exists $\zeta\in(0,1)$ such that $\Psi_{\zeta}(\tilde{x},y):=\zeta|\tilde{x}|^{2}+(1-\zeta)W$ is a FxT ISS Lyapunov function for the $(\tilde{x},y)$ system if $\varepsilon\in(0,\varepsilon^{*})$. Moreover, the second part of Theorem 2 tells us that that function
$\tilde{\Psi}_{\zeta}(\tilde{x},\tilde{z}):=\Psi_{\zeta}(\tilde{x},\tilde{z}-\tilde{x})=\zeta|\tilde{x}|^{2}+(1-\zeta)W(\tilde{z}-\tilde{x})$
is a FxT ISS Lyapunov function for (31) for $\varepsilon\in(0,\varepsilon^{*})$. Since the only difference between (28) and (31) is the perturbation in (28b), we can take the Lie derivative of $\tilde{\Psi}_{\zeta}$ along (28) to obtain:
$\dfrac{d}{dt}\tilde{\Psi}_{\zeta}(\tilde{x},\tilde{z})$ $\leq-c_{1}\tilde{\Psi}_{\zeta}^{\gamma_{1}}-c_{2}\tilde{\Psi}_{\zeta}^{\gamma_{2}}+\rho_{\varepsilon,\zeta}(|\varepsilon_{0}|)$
$~{}~{}~{}+\varepsilon_{0}\frac{\partial W}{\partial y}\bigg{\rvert}_{y=\tilde{z}-\tilde{x}}\dfrac{d\varphi(\theta)}{d\theta}\Pi(\theta)$
$\leq-c_{1}\tilde{\Psi}_{\zeta}^{\gamma_{1}}-c_{2}\tilde{\Psi}_{\zeta}^{\gamma_{2}}+\rho_{\varepsilon,\zeta}(|\varepsilon_{0}|)+M\eta|\varepsilon_{0}|(|\tilde{z}|+|\tilde{x}|)$
$\leq-\frac{c_{1}}{2}\tilde{\Psi}_{\zeta}^{\gamma_{1}}-\frac{c_{2}}{2}\tilde{\Psi}_{\zeta}^{\gamma_{2}}+\rho_{\varepsilon,\zeta}(|\varepsilon_{0}|)+M\overline{\alpha}(c)|\varepsilon_{0}|^{2}$
$~{}~{}~{}-\frac{c_{1}}{2}\tilde{\Psi}_{\zeta}^{\gamma_{1}}-\frac{c_{2}}{2}\tilde{\Psi}_{\zeta}^{\gamma_{2}}+\frac{M}{2\underline{\alpha}(c)}(|\tilde{z}|^{2}+|\tilde{x}|^{2}),$ (34)
where $\overline{\alpha},\underline{\alpha}\in\mathcal{K}_{\infty}$ are obtained from Lemma 1. Now, we make the following observation
$\tilde{\Psi}_{\zeta}(\tilde{x},\tilde{z})$ $\geq|\tilde{x}|^{2}+(1-\zeta)|\tilde{z}|^{2}-2(1-\zeta)|\tilde{z}||\tilde{x}|$
$=\begin{bmatrix}|\tilde{x}|\\ |\tilde{z}|\end{bmatrix}^{\top}\begin{bmatrix}1&\zeta-1\\ \zeta-1&1-\zeta\end{bmatrix}\begin{bmatrix}|\tilde{x}|\\ |\tilde{z}|\end{bmatrix}$
$>\mu\left(|\tilde{x}|^{2}+|\tilde{z}|^{2}\right)$
which holds for $\mu>0$ sufficiently small (this is possible since $\zeta\in(0,1)$). Next, we continue from (34) as follows:
$\dfrac{d}{dt}\tilde{\Psi}_{\zeta}(\tilde{x},\tilde{z})$ $\leq-\frac{c_{1}}{2}\tilde{\Psi}_{\zeta}^{\gamma_{1}}-\frac{c_{2}}{2}\tilde{\Psi}_{\zeta}^{\gamma_{2}}+\rho_{\varepsilon,\zeta}(|\varepsilon_{0}|)+M\overline{\alpha}(c)|\varepsilon_{0}|^{2}$
$~{}~{}~{}-\frac{c_{1}\mu^{\gamma_{1}}}{2}(|\tilde{z}|^{2}+|\tilde{y}|^{2})^{\gamma_{1}}-\frac{c_{2}\mu^{\gamma_{2}}}{2}(|\tilde{z}|^{2}+|\tilde{y}|^{2})^{\gamma_{2}}$
$~{}~{}~{}+\frac{M}{2\underline{\alpha}(c)}(|\tilde{z}|^{2}+|\tilde{x}|^{2}).$
We can then pick $c$ sufficiently large such that
$\frac{M}{2\underline{\alpha}(c)}<\min\left(\frac{c_{1}\mu^{\gamma_{1}}}{2},\frac{c_{2}\mu^{\gamma_{2}}}{2}\right),$
and then apply Lemma 2 to obtain the result.
To further illustrate our results, we simulate system (27) with the cost function $\phi_{\theta}$ taking the quadratic form $\phi_{\theta}(z)=\frac{1}{2}x^{\top}Q_{\theta}z+b_{\theta}^{\top}z$, where $Q_{\theta}$ and $b_{\theta}$ are given by
$Q_{\theta}:=\begin{bmatrix}3+d_{1}&2\\ 2&5+d_{2}\end{bmatrix},\quad b_{\theta}:=\begin{bmatrix}2+d_{3}\\ 1\end{bmatrix}.$
@@
The parameters $d_{i}$ are given by $d_{1}(t)=0.8\sin(2.2\varepsilon_{0}t)$, $d_{2}(t)=1.8\sin(1.7\varepsilon_{0}t)$, $d_{3}(t)=0.66\sin(1.9\varepsilon_{0}t)$. These signals can be generated by a system of the form (24) by setting $\theta\in\mathbb{R}^{6}$, $d_{i}(t)=\theta_{2i}(t)$, and $\Pi(\theta)=\mathcal{R}\theta$, where $\mathcal{R}\in\mathbb{R}^{6\times 6}$ is a block diagonal matrix with rotation matrices on the diagonal. Moreover, in accordance with Proposition 3 we set $\xi_{1}=\frac{1}{3},\xi_{2}=-\frac{1}{5},\eta_{1}=\frac{2}{5},\eta_{2}=-\frac{2}{7}$. It can be verified that $Q_{\theta}\succ 0$ for all $[d_{1},d_{2}]^{\top}\in[-0.8,0.8]\times[-1.8,1.8]$, and the optimizer of $\phi_{\theta}$ is given by $z^{*}=-Q_{\theta}^{-1}b_{\theta}$. As observed in Figure 2, the state $z$ can accurately track the time-varying optimizer in fixed-time.
@@ 5 Conclusion
For this work we establish sufficient interconnections that can guarantee FxT ISS for singularly perturbed systems provided a sufficiently large time scale separation between the two subsystems. Moreover, we show derive conditions on the quasi-steady state map that can allow one to leverage FxT ISS for the error system (7) to obtain FxT ISS for the original system (1). The results were applied to two illustrative examples: a particular nonsmooth second-order interconnection of systems, and a general fixed-time feedback optimization problem with time-varying cost functions, which has not been addressed before using fixed-time stability tools. While these examples showcase the applicability of our results, we hope that our method of verifying the interconnection conditions establishes an efficient paradigm for applying our results to other classes of algorithms and feedback schemes that exhibit multiple time scales. Future directions include applying our results to a broader range of systems, generalizing to systems with more than two time scales, and developing different techniques, such as small gain theory and homogeneity, to study FxTS in interconnected systems.

### v2

@@ 4.1 Fixed-Time Gradient-Based Feedback
$\left\lvert\begin{bmatrix}x(\tau)\\ z(\tau)-h(x(\tau)+\varphi(\theta(\tau)))\end{bmatrix}\right\rvert$
$~~~~~~~~~\leq\beta\left(\left\lvert\begin{bmatrix}x_{0}\\ z_{0}-h(x_{0}+\varphi(\theta_{0}))\end{bmatrix}\right\rvert,\tau\right)+\varrho(|u(\tau)|_{\infty}),$
for all $\varepsilon\in(0,\varepsilon^{*})$, $\tau\geq 0$ and $\varepsilon_{0}\geq 0$. $\square$
Proof 4.1.
Consider the $\tau$-time scale system (36a). This system has a quasi-steady state $h(x+\varphi(\theta))$, which results in the following reduced system:
$\frac{dx}{d\tau}$ $=-\mathcal{F}_{\xi_{1},\xi_{2}}(\nabla\Phi_{\theta}(x+\varphi(\theta)))-\varepsilon_{0}\mathbf{J}_{\varphi}(\theta)\Pi(\theta).$ (37)
[…]
$\frac{dV}{d\tau}\leq-\frac{2}{L}|\nabla\Phi_{\theta}(x+\varphi(\theta))|^{2-\xi_{1}}-\frac{2}{L}|\nabla\Phi_{\theta}(x+\varphi(\theta))|^{2-\xi_{2}}$
$~~~+c|x|^{2}+\frac{M^{2}}{c}|\varepsilon_{0}\Pi(\theta)|^{2}$
$\leq-\frac{\kappa^{2-\xi_{1}}}{L}V^{1-\frac{1}{2}\xi_{1}}{(x)}-\frac{\kappa^{2-\xi_{2}}}{L}V^{1-\frac{1}{2}\xi_{2}}{(x)}+\frac{M^{2}}{c}|\varepsilon_{0}\Pi(\theta)|^{2},$
where $0<c<\frac{1}{L}\min\left\{{\kappa^{2-\xi_{1}}},{\kappa^{2-\xi_{2}}}\right\}$, which shows that Assumption 2 is satisfied. Next, let $y=z-h(x+\varphi(\theta))$, which leads to the following dynamics:
$\frac{dy}{d\tau}$ $=\frac{1}{\varepsilon}g(x+\varphi(\theta),y+h(x+\varphi(\theta)))$ (38)
$~~~-\mathbf{J}_{h}(x+\varphi(\theta))\left(\frac{dx}{d\tau}+\varepsilon_{0}\mathbf{J}_{\varphi}(\theta)\Pi(\theta)\right).$
This yields the following boundary layer system:
$\dot{y}=g(x+\varphi(\theta),y+h(x+\varphi(\theta))).$ (39)
It is easy to verify that Assumption 3 holds with Lyapunov function $W(y)$ obtained from Assumption 4, so it remains to check the interconnection conditions (15a)-(15b). Indeed, if we denote $P_{\theta}(x,y):=\hat{P}_{\theta}(x+\varphi(\theta),y+h(x+\varphi(\theta)))$, we have:
$|I_{1}|$ $\leq 2|x|\bigg{(}\left\lvert\frac{P_{\theta}(x,y)}{|P_{\theta}(x,y)|^{\xi_{1}}}-\frac{P_{\theta}(x,0)}{|P_{\theta}(x,0)|^{\xi_{1}}}\right\rvert$
$~~~+\left\lvert\frac{P_{\theta}(x,y)}{|P_{\theta}(x,y)|^{\xi_{2}}}-\frac{P_{\theta}(x,0)}{|P_{\theta}(x,0)|^{\xi_{2}}}\right\rvert\bigg{)}$
$\leq 2^{\xi_{1}+1}\ell^{1-\xi_{1}}|x||y|^{1-\xi_{1}}$
$~~~+2K|x||y|\left(\ell^{-\xi_{2}}|y|^{-\xi_{2}}+L^{-\xi_{2}}|x|^{-\xi_{2}}\right)$
$\leq c\left(|x|^{2-\xi_{1}}+|x|^{2-\xi_{2}}\right)+\frac{1}{\mu_{1}(c)}\left(|y|^{2-\xi_{1}}+|y|^{2-\xi_{2}}\right)$
$\leq c\tilde{V}^{2}(x)+\frac{c_{1}^{\frac{1}{2}\xi_{1}-1}+c_{1}^{\frac{1}{2}\xi_{2}-1}}{\mu_{1}(c)}\tilde{W}^{2}(y),$
for all $c>0$, and $\mu_{1}\in\mathcal{K}_{\infty}$ is given by $\mu_{1}(c):=m^{-1}\left(c^{-\sigma_{1}}+c^{-\sigma_{2}}+c^{-\frac{1}{\sigma_{2}}}\right)^{-1}$, with $\sigma_{i}=\frac{1}{1-\xi_{i}}$ and $m:=\max\{(2^{\xi_{1}+1}\ell^{1-\xi_{1}})^{1+\sigma_{1}},(4K\ell^{-\xi_{2}})^{1+\sigma_{2}},(4K\ell^{-\xi_{2}})^{1+\frac{1}{\sigma_{2}}}\}$. Picking $c<\frac{1}{L}\min\left\{{\kappa^{2-\xi_{1}}},{\kappa^{2-\xi_{2}}}\right\}$ establishes (15a) and (15c). To verify (15b), we have
$I_{2}$ $=\nabla W(y)^{\top}\mathbf{J}_{h}(x+\varphi(\theta))\mathcal{F}_{\xi_{1},\xi_{2}}(P_{\theta}(x,y)).$ (40)
Let $h:=\sup_{x}|\mathbf{J}_{h}(x)|$, which exists and is finite since $h$ is $\mathcal{C}^{1}$ and globally Lipschitz. We can proceed from (40) as follows:
$|I_{2}|$ $\leq\eta h|y|\left((\ell|y|+L|x|)^{1-\xi_{1}}+(\ell|y|+L|x|)^{1-\xi_{2}}\right)$
$\leq\tilde{L}\left(|x|^{2-\xi_{1}}+|x|^{2-\xi_{2}}+|y|^{2-\xi_{1}}+|y|^{2-\xi_{2}}+|y|^{2}\right)$
$\leq\tilde{L}\left(\tilde{V}^{2}+3\tilde{W}^{2}\right),$
where $\tilde{L}:=\eta h\max\{\ell^{1-\xi_{1}},L^{1-\xi_{1}},2^{-\xi_{2}}\ell^{1-\xi_{2}},2^{-\xi_{2}}L^{1-\xi_{2}}\}$. By applying Corollary 2 we obtain the result.
[…]
$Q_{\theta}:=\begin{bmatrix}3+d_{1}{(t)}&2\\ 2&5+d_{2}{(t)}\end{bmatrix},\quad b_{\theta}:=\begin{bmatrix}2+d_{3}{(t)}\\ 1\end{bmatrix}$.
@@
The parameters $d_{i}$ are given by $d_{1}(t)=0.8\sin(2.2\varepsilon\varepsilon_{0}t)$, $d_{2}(t)=1.8\sin(1.7\varepsilon\varepsilon_{0}t)$, $d_{3}(t)=0.66\sin(1.9\varepsilon\varepsilon_{0}t)$. These signals can be generated by a system of the form (29) by setting $\theta(t)\in\mathbb{R}^{6}$, $d_{i}(t)=\theta_{2i}(t)$, and $\Pi(\theta)=\mathcal{R}\theta$, where $\mathcal{R}\in\mathbb{R}^{6\times 6}$ is a block diagonal matrix with rotation matrices on the diagonal. Moreover, in accordance with Proposition 4 we set $\xi_{1}=\frac{1}{3},\xi_{2}=-\frac{1}{5}$. It can be verified that $Q_{\theta}\succ 0$ for all $t\geq 0$, and the optimizer of $\Phi_{\theta}$ is given by $\varphi(\theta)=-Q_{\theta}^{-1}b_{\theta}$. We interconnect (34b) with the plant dynamics, where $\hat{P}_{\theta}(\hat{x},z)=\frac{1}{2}Q_{\theta}z+Q_{\theta}\hat{x}+2b_{\theta}$. The trajectories of the system are shown in Figure 2, with $\varepsilon=0.05$ and different values of $\varepsilon_{0}$. As observed in the plot, the state $z$ converges in fixed-time to a neighborhood of the time-varying optimizer, whose size shrinks as $\varepsilon_{0}\to 0^{+}$.
@@ 5 Conclusion
We establish sufficient Lyapunov conditions for the study of FxT ISS properties in singularly perturbed systems. The results were applied to two illustrative examples: a particular nonsmooth second-order interconnection of systems, and a general fixed-time feedback optimization problem with time-varying cost functions, which has not been addressed before using fixed-time stability tools. Our method of verifying the interconnection conditions establishes an efficient paradigm for applying our results to other classes of algorithms and feedback schemes that exhibit multiple time scales. Future research directions include applying our results to a broader range of systems, including systems with more than two time scales. Moreover, it is also of interest to identify a more general class of systems and quasi-steady state mappings for which our interconnection conditions hold, including characterizations based on homogeneity.